# Twisted Superfields

### v7

@@ 5.1 A Light Z’ Boson
At the electroweak scale, the generalized Weinberg rotation of eq (2.16) rotates the $T_{2}^{3}$, $T^{Y}$, and $T^{Y^{\prime}}$ matrices of eqs (1.7) and (2.15) to $T^{\gamma}$, $T^{Z}$, and $T^{Z^{\prime}}$. Using the angles from eq (2.17), the $Z^{\prime}$ matrices take the form:
$g_{Z^{\prime}}T_{1}^{Z^{\prime}}$ $=-{\rm{diag}}\left(1+\tfrac{2}{3}x,1+\tfrac{2}{3}x,1+\tfrac{2}{3}x,0,x-1,x-1\right)e\sin\phi_{Z}/\sin\theta_{W}\cos\theta_{W}$ (5.1)
$g_{Z^{\prime}}T_{2}^{Z^{\prime}}$ $=-{\rm{diag}}\left(\tfrac{1}{3}-\tfrac{1}{3}x,\tfrac{1}{3}-\tfrac{1}{3}x,\tfrac{1}{3}-\tfrac{1}{3}x,\tfrac{1}{6}-x,-\tfrac{5}{6},-\tfrac{5}{6}\right)e\sin\phi_{Z}/\sin\theta_{W}\cos\theta_{W}$
$g_{Z^{\prime}}T_{3}^{Z^{\prime}}$ $={\rm{diag}}\left(\tfrac{1}{2}+\tfrac{1}{3}x,\tfrac{1}{2}+\tfrac{1}{3}x,\tfrac{1}{2}+\tfrac{1}{3}x,x-1,0,0\right)e\sin\phi_{Z}/\sin\theta_{W}\cos\theta_{W}$
$g_{Z^{\prime}}T_{A}^{Z^{\prime}}$ $={\rm{diag}}\left(\tfrac{31}{12}+\tfrac{1}{3}x,\tfrac{31}{12}+\tfrac{1}{3}x,\tfrac{31}{12}+\tfrac{1}{3}x,-\tfrac{37}{12}+x,-\tfrac{25}{12},-\tfrac{25}{12}\right)e\sin\phi_{Z}/\sin\theta_{W}\cos\theta_{W}\,,$
where the last line is the group structure of $Z^{\prime}$ acting on adjoint fields, and $x=\sin^{2}\theta_{Z}\simeq\sin^{2}\theta_{W}$ for small $\sin\phi_{Z}$. The group structure of the $Z^{\prime}$ acting on gauginos is the same as that of $T_{2}^{Z^{\prime}}$.
Given those group structures and the definitions of quarks and leptons earlier in the paper, the following couplings with the $Z^{\prime}$ boson can be read off of eq (5.1):
$\begin{array}[]{|r|c|c|}\hline\cr&g^{L}&g^{R}\\ \hline\cr u&-1-\tfrac{2}{3}x&-\tfrac{1}{6}-\tfrac{2}{3}x\\ d&\tfrac{14}{3}+\tfrac{1}{3}x&-\tfrac{7}{6}+\tfrac{1}{3}x\\ e&1-x&1-x\\ \nu_{e}&0&0\\ \hline\cr\end{array}\,\,\,\times\frac{e\sin\phi_{Z}}{\sin\theta_{W}\cos\theta_{W}}$ (5.2)
@@
$\begin{array}[]{|r|c|c|}\hline\cr&V&A\\ \hline\cr p&C_{p}\simeq 0.12&a_{p}\simeq 0.49\\ n&C_{n}\simeq 1.0&a_{n}\simeq-0.95\\ e&|\epsilon|\simeq 0.87&0\\ \hline\cr\end{array}\,\,\,\times 10^{-3}\,,$ (5.4)
@@
The mass of the $X17$ in this model is generated by its interaction with the scalar field $\tilde{\phi}_{22}$, and its vev $\bar{\tilde{\phi}}_{22}$. From the form of $g_{Z^{\prime}}T_{2}^{Z^{\prime}}$ in eq (5.1), the tree-level mass is given by:
$m^{2}_{Z^{\prime}}\simeq\bar{\tilde{\phi}}_{22}^{2}\left(\tfrac{5}{6}\right)^{2}\left(\frac{e\sin\phi_{Z}}{\sin\theta_{Z}\cos\theta_{Z}}\right)^{2}\,.$ (5.5)
From the quark mass matrices, the vev $\bar{\tilde{\phi}}_{22}$ should be only a couple of GeV. If so, then the tree-level $Z^{\prime}$ mass is only about 2 MeV. It is assumed that the remainder of its 17 MeV mass comes from radiative corrections.
5.2 A Seventh Quark
This model predicts the existence of a seventh flavor of quark. The model itself does not predict the mass of the quark, but if the mass was larger than that of the top quark, then the seventh quark could potentially break electroweak symmetry at its mass scale, not that of the top quark. In other words, the model appears to prefer a seventh quark with a mass smaller than that of the top quark.
A recent paper shows that if there is a fourth down-type quark with a mass of 2.8 GeV, then most of the exotic hadrons discovered over the last twenty years fit nicely into the quark model as normal mesons and baryons involving this quark, rather than 4- or 5-quark hadrons [40].
If a fourth down-type quark existed, one might expect it to generate predictions for CKM data that no longer agree with experimental data. The opposite is true. The CKM matrices of eq (3.22) accurately reproduce absolute values of CKM element data, even for some data that disagree with the Standard Model by $\sim 3\sigma$. In particular, the model reproduces data that cast doubt on CKM unitarity and on the W boson only connecting left-handed quarks [41].
A natural question is how an additional 2.8 GeV quark could have evaded detection so far. Most of the dozens of direct searches for a new down-type quark surprisingly do not rule out a quark of the type proposed here [42]. In particular, most model-independent searches do not exclude a new quark with a mass smaller than that of the beauty quark.
One notable exception is the fact that inclusive hadronic cross section data seem to rule out the possibility of a light seventh quark. The Standard Model reproduces these data very well, apparently leaving no room for an additional quark. However, it was recently proposed that an assumption inherent in Standard Model calculations may be incorrect [43]. Namely, calculations assume that there is no interference between $Z/\gamma^{*}\to u\bar{u}\to X$ and $Z/\gamma^{*}\to d\bar{d}\to X$ events, where $X$ is the same hadronic state in both processes (e.g. $\pi^{+}\pi^{-}$). If interference is taken into account, the Standard Model may underpredict inclusive hadronic cross section data below the $Z$ pole. If so, the additional quark may be required in order to fit the experimental data.
Another difficulty with the model is the fact that the charm quark of this model has a different coupling to the $Z$ boson than in the Standard Model. In [41], a mechanism is proposed whereby the effects of the different charm coupling combined with the additional quark decaying to charm allow this model to reproduce the LEP-measured charm cross section and asymmetry data.
But the mechanism described there would not help reconciling $Z$ boson width measurements with the larger total coupling of the charm quark in this model and the presence of an additional quark. A mechanism that could allow this model to reproduce those data is described in the $Z$ width section below.
Under the hypothesis that the seventh quark exists and has a mass of around 2.8 GeV, a large number of predictions have been made for new exotic hadrons that could be observed in current experiments [40].
5.3 Right-Handed Neutrinos
In the neutrino part of section 3, it was proposed that $\nu_{W}^{(2)}$ and $\nu_{W}^{(3)}$ are right-chiral neutrinos that are isodoublets with the right-chiral muon and tau lepton, respectively. These right-chiral neutrinos have Dirac mass terms together with two of the model’s isosinglet neutrinos ($\tilde{\nu}^{(2)}$ and $\tilde{\nu}^{(3)}$). That makes the heavy right-chiral neutrinos of this model different than most models that form the basis for experimental searches. In those models, the heavy right-chiral neutrino is usually an isosinglet that has a Dirac mass term mixing it with one of the observed left-chiral isodoublet neutrinos. This model difference means that some of the heavy neutrino searches do not set limits on this model.
With that background, it is assumed that the masses of $\nu_{W}^{(2)}$ and $\nu_{W}^{(3)}$ are an MeV or so heavier than the mass of the tau lepton. Despite many searches that produced no evidence for heavy neutral leptons, it appears that neutrinos of the masses and type proposed here may not be ruled out.
Figure 4 of [44] shows exclusion limits for a massive neutrino that interacts with the tau lepton. The only existing (as opposed to projected) experiment shown in that figure pertinent to the mass of the heavy tau neutrino proposed above is DELPHI. The DELPHI experiment searched for $Z\to\nu_{h}\bar{\nu}$ events, where $\nu_{h}$ is a heavy neutrino and $\bar{\nu}$ is a normal (almost massless) antineutrino [45]. In the model proposed here, the $Z$ boson has no direct coupling of the heavy right-chiral $\nu_{W}^{(3)}$ tau neutrino with the normal neutrinos $\nu_{W}^{(1)}$, $\tilde{\nu}_{W}^{(2)}$ and $\tilde{\nu}_{W}^{(3)}$. Also, the model’s mass parameters can accommodate heavy-light neutrino mass mixing that is below the DELPHI limit. In the simplified example matrix of eq (3.27), that would require choosing parameters such that $m_{133}<.01m_{\nu 3}$. In this way, the model can evade the DELPHI limits for a heavy tau neutrino.
Figure 1 of [46] shows exclusion limits for a massive neutrino that interacts with the muon. Experiments performed by DELPHI, L3 [47], CHARM II [48] and NA3 [49] limit the mass. L3 looked for heavy neutrinos in the same way as DELPHI (searching for $Z\to\nu_{h}\bar{\nu}$ events). CHARM II searched for transitions from normal muon neutrinos to heavy ones in neutral current scattering. For $m_{122}<.005m_{\nu 2}$, the right-chiral muon neutrino of this model evades the DELPHI, L3 and CHARM II limits. NA3 was not sensitive to a heavy neutrino as short-lived as the one considered here. So the heavy right-chiral muon neutrino considered here is not excluded by the experiments quoted in [46].
A CMS search for heavy neutrinos from $W$ decay only considered squared interaction strengths up to 0.5 [50]. Since the neutrino considered here has an interaction strength of 1, it is not excluded.
Despite its short life, the CMS search from $B$ meson decay [51] could potentially have seen some signal of the heavy muon neutrino considered here. That experiment ruled out a neutrino of this type, unless its mass was in their veto range of 1.74 to 1.8 GeV. The reason for the veto is because that range could be contaminated by an unknown number of $D^{0}\to K\pi$ decays. The CMS data did show a feature in that range, but it was attributed to $D^{0}$ decays and excluded from the analysis. It is assumed that some part of that feature was due to the heavy muon neutrino proposed here.
Neutrinos of the type discussed here are too heavy to be produced in the decay of the $D$ meson or tau lepton, so their presence would not contradict $D$ or tau decay branching ratios or the fact that measured Michel parameters are consistent with only left-chiral leptons being produced in decays of tau leptons [52, 53].
The heavy neutrinos proposed here are not sterile; they interact with the $Z$ boson. Therefore, they add to the inclusive cross sections that were measured very precisely near the $Z$ peak at LEP. A way to reconcile this model with those measurements is discussed in the $Z$ width section below.
As noted in [5], experimentally observed branching ratios for $B\to D^{(*)}\tau\nu$ exceed Standard Model predictions by $>3\sigma$. Also, the number of observed multilepton LHC events exceeds Standard Model predictions by $>5\sigma$. This model provides the following new physics explanation for those anomalies.
Heavy neutrinos of the type proposed here could be produced in the decay $B\to D^{(*)}\mu\nu_{W}^{(2)}$ with $\nu_{W}^{(2)}$ then decaying to two charged particles (accompanied by a normal neutrino for pure leptonic decays). Since the mass of $\nu_{W}^{(2)}$ is similar to that of the tau lepton, these decays could look like the decay $B\to D^{(*)}\tau\nu$, with $\tau$ decaying to three charged particles. If an experiment grouped these decays that way (or grouped some of them that way), then the experimental results would show an apparent branching ratio for $B\to D^{(*)}\tau\nu$ decays that was significantly larger than the branching ratio predicted by the Standard Model. It is proposed that this is the origin of the $B\to D^{(*)}\tau\nu$ anomaly.
In LHC events, the heavy neutrinos of this model lead to many more multilepton events. As an example, it is interesting to consider a $t\bar{t}$ event, where the $t$ decays to a b-jet and a $W^{+}$ boson. In the Standard Model, there is some probability that the $W^{+}$ produces a charged lepton and a neutrino. In this model, that probability is doubled for production of muons and tau leptons, due to production involving their heavy neutrinos. Those heavy neutrinos also decay to at least one additional charged lepton. The net result is that this model predicts many more multilepton events at LHC than does the Standard Model.
5.4 Muon and Tau Asymmetry Data
The muon and tau lepton of this model have different couplings to the Z boson than in the Standard Model. Those differences must be reconciled with the fact that cross section and asymmetry data from $e^{+}e^{-}\to Z/\gamma^{*}\to\mu^{+}\mu^{-}$ or $\tau^{+}\tau^{-}$ decays are consistent with the muon and tau lepton having the same coupling to the Z boson as in the Standard Model [54].
The couplings to the Z boson for the left- and right-chiral muons and tau leptons for the Standard Model (SM) and this model (Model) are:
$\begin{array}[]{|r|c|c|}\hline\cr\textrm{Z coupling}&g^{L}_{\mu,\tau}&g^{R}_{\mu,\tau}\\ \hline\cr\mu,\tau\textrm{ (SM)}&-\tfrac{1}{2}+\sin^{2}\theta_{W}&\sin^{2}\theta_{W}\\ \mu,\tau\textrm{ (Model)}&-\tfrac{1}{2}+\sin^{2}\theta_{W}&-\tfrac{1}{2}+\sin^{2}\theta_{W}\\ \hline\cr\end{array}\,.$ (5.6)
It can be seen that if some of this model’s new physics was able to modify the effective right-chiral Z-couplings of the muon and tau lepton by $\delta g^{R}_{\mu,\tau}=\tfrac{1}{2}$, then their effective Z-couplings would be the same as those in the Standard Model, and consistent with the data.
From eqs (2.18) and (2.20), it can be seen that the right-handed muon and tau neutrinos discussed in the last section ($\nu_{W}^{(2)}$ and $\nu_{W}^{(3)}$) have couplings with the Z boson of $\tfrac{1}{2}$.
In this model, there would have been events at LEP where a $Z$ boson decayed to $\nu_{W}^{(3)}\nu_{W}^{(3){\dagger}}$. If the mass of $\nu_{W}^{(3)}$ is only slightly larger than the tau lepton, then it will decay by $\nu_{W}^{(3)}\to\tau^{-}e^{+}\nu_{e}$. Therefore, $Z\to\nu_{W}^{(3)}\nu_{W}^{(3){\dagger}}$ events would have looked like $Z\to\tau^{-}\tau^{+}e^{+}e^{-}$ events. If the experiments considered $e^{+}e^{-}$ to be radiated pairs, then these events would have been grouped together with the remaining $Z\to\tau^{-}\tau^{+}$ events. In that case, the effective right-chiral tau lepton coupling produced by this model would have been very similar to that of the Standard Model and therefore would have reproduced the data.
Similarly, decays of $Z\to\nu_{W}^{(2)}\nu_{W}^{(2){\dagger}}$ may have been grouped together with $Z\to\mu^{-}\mu^{+}$ events. Again, the effective right-chiral muon coupling of this model would then have been very similar to that of the Standard Model.
There is a subtlety with this scenario. If 4-lepton events had been rejected by the tagging at LEP or SLD, then a significant number of right-chiral neutrino events would have been excluded from effective coupling data. However, it appears that these events were not excluded. It was stated that “The classification of final states with radiated fermion pairs, i.e. of four-fermion events, into one of the three lepton categories is made by choosing the lepton pair with the highest invariant mass” [54].
It is also interesting that about 1.2% of $e^{+}e^{-}\to\mu^{+}\mu^{-}$ events observed by OPAL at LEP had two muons with the same charge. It was mentioned that “the origin of this effect is not understood” [55], and it was assumed that one of the muons had been misidentified. An alternative explanation in this model is provided by events involving the decay $\nu_{W}^{(2)}\to\mu^{-}\mu^{+}\nu_{\mu}$. Rather than pair the $\mu^{-}$ from $\nu_{W}^{(2)}$ with the $\mu^{+}$ from $\nu_{W}^{(2){\dagger}}$, some events could have paired the $\mu^{+}$ from $\nu_{W}^{(2)}$ with the $\mu^{+}$ from $\nu_{W}^{(2){\dagger}}$.
In summary, it is proposed that the decay of Z bosons to right-chiral neutrinos that were then tagged as right-chiral muons and tau leptons could have allowed the model presented here to have reproduced the muon and tau lepton cross section and asymmetry measurements at LEP and SLD.
5.5 The Z boson width
Precision experiments performed at LEP and SLD established the width of the $Z$ boson and how much of that width was due to the $Z$ boson decaying to hadrons, charged leptons, and invisible particles (light neutrinos) [54]. The model presented here has additional Beyond-Standard-Model particles that would have also been produced in $Z$ boson decay: an additional down-type quark and two additional neutrinos with masses similar to that of the tau lepton. The model also has $Z$ interactions with the right-chiral muon, tau lepton and charm quark that are larger than in the Standard Model. This subsection qualitatively presents a way that this model could potentially reproduce the $Z$ boson width data.
The LEP hadronic cross sections were calculated by counting hadronic events (less background) and dividing by the luminosity (and efficiency). Leptonic cross sections were calculated in a similar way. If it is assumed that the actual luminosity of these events was $\sim$13lower than the luminosity used in the analysis of the LEP experiments, then it would be possible for this model to reproduce the LEP $Z$ boson width data (along with the various ratios of hadronic cross sections to leptonic cross sections).
At LEP, the luminosity was estimated by measuring the number of low angle $e^{+}e^{-}\to e^{+}e^{-}$ events (angles of $\sim 25-100$ milliradians relative to the beam line). If the Standard Model is assumed, then these events are almost entirely due to t-channel photon exchange Bhabha scattering. Since only QED is involved, this process can be calculated with high accuracy, leading to very precise luminosity estimates (see for example [56]).
In this model, there are other processes that contribute to low angle $e^{+}e^{-}\to e^{+}e^{-}$ events. First, a t-channel diagram involving the $Z^{\prime}$ boson of this model can contribute, and its diagram interferes constructively with the photon exchange diagram. Approximating the $Z^{\prime}$ boson as massless, from eq (5.1), the t-channel Bhabha scattering cross section in this model compared to that of the Standard Model is:
$\frac{\sigma_{Bh}}{\sigma^{SM}_{Bh}}\simeq\left(1+\frac{\sin^{2}\phi_{Z}}{\tan^{2}\theta_{W}}\right)^{2}\,.$ (5.7)
A value of $\sin\phi_{Z}\simeq 0.1$ at the $Z$ boson mass scale would translate into about a 7% increase in t-channel $e^{+}e^{-}\to e^{+}e^{-}$ events, and a luminosity about 6% smaller than calculated using the Standard Model.
Another process could potentially have an even larger impact. In section 3, it was proposed that the electron’s mass is generated by its interaction with the adjoint scalar $(\varphi_{2})_{32}$ that has a vev of $\bar{\varphi}_{2}$. As discussed in the scalar section below, the mass of this scalar could be small, potentially even smaller than that of the electron.
If $(\varphi_{2})_{32}$ does have a small mass, the t-channel exchange of this scalar could also generate more low angle $e^{+}e^{-}\to e^{+}e^{-}$ events for the same beam luminosity than the Standard-Model prediction. Between the $Z^{\prime}$ and this scalar, the model could produce $\sim$15more t-channel $e^{+}e^{-}\to e^{+}e^{-}$ events than the Standard Model. That would imply that a luminosity calculated using this model would be $\sim$13lower than the luminosity calculated using the Standard Model.
If that were the case, then the number of hadronic events observed would correspond to a $\sim$15larger hadronic cross section than predicted by Standard-Model calculations, so Standard-Model-only $Z$ boson decays would not be able to reproduce the data. The same is true of leptonic decays. The model presented here, however, could reproduce the observed cross sections due to the additional particles and different $Z$ couplings mentioned at the beginning of this section.
The $(\varphi_{2})_{32}$ scalar also interacts with the $Z$ boson, so it would be produced in pairs in $Z$ boson decays. Due to its low mass and zero charge, $(\varphi_{2})_{32}$ would add to the invisible width of the $Z$ boson.
But this scalar would also contribute to the initial state radiation that significantly increases the actual width of the $Z$ boson compared to what is observed (see for example fig. 1.12 of [54]). That additional increase in width could potentially accommodate the additional invisible partial width of $(\varphi_{2})_{32}$.
To end this section, it is worth pointing out that the example value of $\sin\phi_{Z}$ mentioned above corresponds to $\cos\phi_{Z}\simeq 0.995$. From eq (3.6), and given the extraordinary precision in measurements of $\sin^{2}\theta_{W}$, such a value would require significant new physics quantum contributions to the $W$ boson mass. This model certainly has a lot of new physics that could make those contributions, and their effect would need to be checked to see if a consistent picture can be achieved.
An alternative scenario is that $\sin\phi_{Z}$ is much smaller at the electroweak scale, so no (or very small) new physics contributions would be needed to reproduce precision $W$ mass measurements. In that scenario, the entire luminosity correction described above is attributed to the $(\varphi_{2})_{32}$ scalar.
5.6 The Muon Magnetic Moment
The main source of uncertainty in a Standard-Model calculation of the $g-2$ magnetic moment of the muon comes from calculation of loop diagrams that involve hadrons. There are two main competing ways to calculate the hadronic contribution: (i) a data-driven approach [57, 58, 59] and (ii) a lattice QCD approach [60].
This model has both an additional quark and also quark couplings that are different from those in the Standard Model. Therefore, a lattice calculation using this model would produce a different result than one using the Standard Model. On the other hand, a data-driven approach based on this model should produce similar results to one based on the Standard Model. The normal data-driven approach assumes that all Standard-Model nonperturbative QCD interactions are incorporated into the low energy data used as input. Changing the model just means changing that assumption (replacing Standard-Model with this model). But since the data are still the same, the hadronic contribution to the calculation would remain the same.
As a result, since no lattice calculations involving this model are available, this model’s $g-2$ results should be compared to those of the data-driven calculation. Also, any differences from the Standard Model result should have their origin in non-hadronic parts of the calculation.
Experimental measurement of the anomalous magnetic moment of the muon produces results that differ from data-driven Standard Model (SM) calculations by 5.1 standard deviations [57, 61, 59]:
$a_{\mu}^{\rm{exp}}-a_{\mu}^{\rm{SM}}=249\pm 48\times 10^{-11}.$ (5.8)
This model may be able to reproduce that value as a result of the model’s differences in the weak sector (as opposed to the hadronic sector).
In the Standard Model, the one-loop contribution to the muon magnetic moment involving the W and Z bosons is [58, 62, 63]:
$a_{\mu}^{\rm{EW}}[\rm{1loop}]=\frac{G_{F}m_{\mu}^{2}}{8\sqrt{2}\pi^{2}}\left(\tfrac{7}{3}+1+\tfrac{4}{3}g_{\mu V}^{2}-\tfrac{20}{3}g_{\mu A}^{2}\right).$ (5.9)
The four terms are from diagrams involving (i) two $W$ bosons, (ii) one $W$ boson and one Goldstone-boson component of the Higgs doublet (fig. 100c of [58]), (iii) a vector current $Z$ boson connection, and (iv) an axial vector $Z$ boson connection. Plugging in the Standard Model couplings, one gets $a_{\mu}^{\rm{EW}}[\rm{1loop}]\simeq 194.8\times 10^{-11}$ [62].
This model’s muon only has a vector-current connection to the $Z$ boson, and it is larger than in the Standard Model. Consequently, the last term of eq (5.9) vanishes, and the third term is larger than in the Standard Model. This model’s muon does not have a tree-level coupling with the Higgs doublet (since the muon does not get its mass from the Higgs boson), so the second term vanishes (or is very small). On the other hand, the right-chiral muon neutrino of this model causes the first term above to get doubled. After plugging in the couplings from eq (5.6), this model generates $a_{\mu}^{\rm{EW}}[\rm{1loop}]\simeq 549\times 10^{-11}$
In the Standard Model, the 2-loop calculation reduces 194.8 down to 153.6 [62], and the 3-loop contribution is negligible. In other words, consideration of 3 loops reduces the 1-loop result by 21%.
Making the very rough assumption that a 3-loop calculation in this model would also reduce the 1-loop result by around 21%, this model’s 3-loop value would be in the range of $433\times 10^{-11}$. If that is compared to the Standard Model value of 153.6, the result is
$a_{\mu}^{\rm{model}}-a_{\mu}^{\rm{SM}}\sim 279\times 10^{-11}.$ (5.10)

### v8

@@ 5.1 A Light Z’ Boson
In section 4, it is mentioned that the top quark generates an anomaly for $Z^{\prime}$. This is actually assumed to first affect $Y^{\prime}$ from the top quark scale to the $Z$ boson mass scale, and then to affect $Z^{\prime}$ below the $Z$ boson mass scale. That anomaly causes the coupling for $Y^{\prime}$ interactions with $Q_{21}$ and gaugino fields (both of which have top-quark components) to run down more quickly than the coupling for $Y^{\prime}$ interactions with $Q_{22}$ fields (which only have bottom and down quark components). That effect can be parametrized by a factor $\kappa_{22}$ by which the $g_{Y^{\prime}}$ coupling to $Q_{22}$ exceeds that to $Q_{21}$ and gaugino fields at the scale of $m_{Z}$.
In that case, then the $T^{Z^{\prime}}_{22}$ group structure is derived by the following modification of the Weinberg rotation of eq (2.16):
$g_{Z^{\prime}}T^{Z^{\prime}}_{22}=$ $\kappa_{22}\cos\phi_{Z}g_{Z^{\prime}}T^{Z^{\prime}}_{22}$ (5.1)
$+$ $\sin\phi_{Z}\left(\cos\theta_{Z}g_{2}T_{2}^{3}-\sin\theta_{Z}g_{Y}T^{Y}_{22}\right)\,.$
Using the angles from eq (2.17), the $Z^{\prime}$ matrices at the 17 MeV scale take the form:
$g_{Z^{\prime}}T_{21}^{Z^{\prime}}=g_{Z^{\prime}}T_{1F}^{Z^{\prime}}$ $=-{\rm{diag}}\left(1+\tfrac{2}{3}x,1+\tfrac{2}{3}x,1+\tfrac{2}{3}x,0,x-1,x-1\right)e\sin\phi_{Z}/\sin\theta_{W}\cos\theta_{W}$ (5.2)
$g_{Z^{\prime}}T_{G}^{Z^{\prime}}$ $={\rm{diag}}\left(\tfrac{1}{2}+\tfrac{1}{3}x,\tfrac{1}{2}+\tfrac{1}{3}x,\tfrac{1}{2}+\tfrac{1}{3}x,x-1,0,0\right)e\sin\phi_{Z}/\sin\theta_{W}\cos\theta_{W}$
$g_{Z^{\prime}}T_{22}^{Z^{\prime}}$ $=\tfrac{1}{2}\kappa^{\prime}_{22}\left(\kappa_{22}{\rm{diag}}(1,1,1,-1,-1,0)+\left(\tfrac{2}{3}x,\tfrac{2}{3}x,\tfrac{2}{3}x,2x-1,1,0\right)\right)e\sin\phi_{Z}/\sin\theta_{W}\cos\theta_{W}\,,$
where $T_{G}^{Z^{\prime}}$ is the group structure when acting on gauginos, and $\kappa^{\prime}_{22}$ is a second parameter that expresses the fact that below $m_{Z}$ but above $m_{b}$, the $g_{Z^{\prime}}$ coupling to $Q_{22}$ runs down more slowly than the $g_{Z^{\prime}}$ coupling to $Q_{21}$ and gaugino fields. The $Z^{\prime}$ group structure when interacting with adjoint fields is assumed to be the same as $T_{22}^{Z^{\prime}}$, since adjoint fields also do not have any top-quark components.
Given those group structures along with the definitions of quarks and leptons earlier in the paper and the quark mass matrices of section 3, $Z^{\prime}$ couplings to left- and right-handed up and down quarks, electrons and neutrinos are:
$\begin{array}[]{|r|c|c|}\hline\cr&g_{L}&g_{R}\\ \hline\cr u&-1.15&-0.65\\ d&3.0&0.71\\ e&-0.15&-0.15\\ \nu_{e}&0&0\\ \hline\cr\end{array}\,\,\,\times\frac{e\sin\phi_{Z}}{\sin\theta_{W}\cos\theta_{W}}\,.$ (5.3)
It was assumed for the fit that $\tfrac{1}{2}\kappa^{\prime}_{22}\left(\kappa_{22}+\tfrac{2}{3}x\right)\simeq 3.0$.
@@
$\begin{array}[]{|r|c|c|}\hline\cr&V&A\\ \hline\cr p&C_{p}\simeq 0.02&a_{p}\simeq 0.25\\ n&C_{n}\simeq 1.1&a_{n}\simeq-0.44\\ e&|\epsilon|\simeq 1.0&0\\ \hline\cr\end{array}\,\,\,\times 10^{-3}\,,$ (5.5)
@@
With these choices, the $Z^{\prime}$ boson of this model can reproduce $X17$ data while satisfying all applicable constraints. From fig. 5 of [33], it can be seen that the proton and neutron vector couplings can reproduce the $X17$ 12C data, while the axial vector couplings can reproduce the 8Be and 4He data. The vector coupling is protophobic enough to satisfy the NA48 constraint. The fact that there are no interactions with electron neutrinos allows the model to satisfy neutrino constraints [34]. The electron interaction strength is above the limit imposed by NA64 but below the level that would contradict electron $g-2$ data [35]. Finally, as described in Appendix F of [33], the very stringent atomic parity constraints on mixed-parity models do not apply when the electron has only a vector interaction (no axial vector), as in this model.
The mass of the $X17$ in this model is generated by the vevs $\bar{\phi}_{22}$ and $\bar{\varphi}_{2}$. From the form of $g_{Z^{\prime}}T_{22}^{Z^{\prime}}$ in eq (5.2) and the expression $m_{b}=1.5g_{2}\bar{\phi}_{22}$ (from eqs (3.15) and (3.14)), the $\bar{\phi}_{22}$ part of the tree-level mass is:
$m_{Z^{\prime}}\simeq m_{b}\left(\tfrac{1}{2}\kappa^{\prime}_{22}(\kappa_{22}-1)\right)\left(\frac{1.5\sin\phi_{Z}}{\cos\theta_{Z}}\right)+...\,.$ (5.6)
This tree-level value from $\bar{\phi}_{22}$ can be as much as 5 MeV, or possibly less, depending on the individual values of $\kappa_{22}$ and $\kappa^{\prime}_{22}$. The remainder of the 17 MeV mass of the $Z^{\prime}$ in this model comes from its interaction with the adjoint vev $\bar{\varphi}_{2}$ and radiative corrections.
5.2 A Seventh Quark
This model predicts the existence of a seventh flavor of quark (denoted here by $f$). The model does not predict the mass of the quark, and by choosing a large value for the superpotential parameter $\Gamma_{222}$, it is possible to construct a version of this model where the $f$ quark has a mass larger than that of the top quark. However, for reasons discussed below, it is proposed that the mass of the $f$ quark is smaller than the mass of the bottom quark.
A natural question is how an additional low-mass quark could have evaded detection so far. A recent paper shows that if there is a fourth down-type quark with a mass of 2.9 GeV, then most of the exotic hadrons discovered over the last twenty years fit nicely into the quark model as normal mesons and baryons rather than as 4- or 5-quark hadrons [36]. In other words, as opposed to evading detection, it is proposed that the additional quark has been observed hundreds of times. In this model, the production and decay processes of hadrons involving the additional quark are mediated by adjoint scalar fields.
In the Scalars subsection below, it is suggested that the scalar $\varphi_{5}$ proportional to diag(1,1,1,1,1,1)111111(1,1,1,1,1,-1)( 1 , 1 , 1 , 1 , 1 , - 1 ) may be very light. This scalar does not interact with any of the gauge bosons or gaugino quarks below the unification scale, but it does interact with quarks in the adjoint twisted superfield. The interaction with those quarks is governed by the superpotential coupling $\Gamma_{\Phi}$ from eq (1.15).
In particular, the model includes the following interactions of quarks with $\varphi_{5}$:
$\Gamma_{\Phi}\varphi_{5}\left(\tilde{u}_{W}^{A}u_{W}^{A}+\tilde{d}_{W}^{A}d_{W}^{A}\right)+h.c.$ (5.7)
$\simeq$ $\Gamma_{\Phi}\varphi_{5}\left(\bar{c}_{L}c_{R}+\left(.22\bar{d}_{L}+.97\bar{s}_{L}+.05\bar{b}_{L}-.04\bar{f}_{L}\right)f_{R}\right)$
$\hskip 14.22636pt+...+h.c.\,,$
where the second line comes from diagonalizing the mass matrices of section 3, and “…” in the last line includes much smaller interactions for the other quark combinations. Eq (5.7) shows that the $\varphi_{5}$ scalar interacts primarily with $c\bar{c}$ as well as with $f\bar{s}$, $f\bar{d}$ and their Hermitian conjugates.
For that reason, one way to produce $f\bar{d}$ and $f\bar{s}$ mesons is via $e^{+}e^{-}\to c\bar{c}\to f\bar{d}$ or $f\bar{s}$, where the first process is mediated by a photon and the second by $\varphi_{5}$. Another production mechanism is via decay of a $b$ quark. For example: $\bar{u}b\to\bar{u}c\bar{c}s\to\bar{u}s+f\bar{d}$, where the first process is mediated by a $W$ boson and the second by $\varphi_{5}$.
The same scalar also mediates the decay of $f$-quark mesons. For example, $\varphi_{5}$ mediates the decay $f\bar{u}\to d\bar{u}+c\bar{c}$. Production and decay mechanisms like these are discussed in detail in [36], enabling this model to reproduce properties of most of the observed exotic hadrons.
An important feature of the model is that the $\bar{b}s$ and $\bar{b}d$ interactions with $\varphi_{5}$ are very small (<.$<.002$). To be consistent with observations, they must be small enough so that scalar-mediated decays such as $b\to sc\bar{c}$ are rarer than weak decays for $B$ mesons and $\Lambda_{b}$. The strength of $\bar{b}f$ is not a problem since $B^{+}$ does not have enough mass to decay to an $\bar{f}u$ meson along with a $c\bar{c}$ or $f\bar{d}$ meson.
The strength of scalar interactions affecting each down-type quark is highly dependent on the details of the model. Eq (5.7) only shows scalar interactions mediated by $\varphi_{5}$. In the next subsection, other scalars are proposed to play a role in certain processes, but in this version of the model, they do not contribute significantly to hadron decay. Another model dependence is the mass matrices of eqs (3.16) and (3.17). There is flexibility to rotate, transform, or otherwise alter those matrices in ways that still reproduce the quark masses, CKM data, and $Z^{\prime}$ data well. This flexibility can be used for example to modify the relative coupling of $\bar{s}f$ vs. $\bar{d}f$ or other details of the model.
If the proposed additional quark exists, one might expect it to generate predictions for CKM data that no longer agree with experimental data. The opposite is true.
@@ 5.2 A Seventh Quark
Vector Current: $|V_{us}|=0.2231\pm 0.0006$ (5.8)
Axial Vector Current: $|V_{us}|=0.2254\pm 0.0006,$